To show that sin^2x+cos^2x=1 is equivalent to tan^2x+1=sec^2x,

what do you need divide sin^2x+cos^2x=1 by?
1
Expert's answer
2020-11-17T17:03:05-0500

Let us divide both parts of the equality "\sin^2x+\cos^2x=1" by "\cos^2 x". Then we have "\frac{\sin^2x+\cos^2x}{\cos^2 x}=\frac{1}{\cos^2{x}}" which is equivalent to "\frac{\sin^2x}{\cos^2 x}+\frac{\cos^2x}{\cos^2 x}=(\frac{1}{\cos{x}})^2" and to "\tan^2x+1=\sec^2x."